OpenStudy (anonymous):

Can someone please help me with trig formulas? I have a big test tomorrow. I will FAN and MEDAL

OpenStudy (freckles):

Do you know what you need help on exactly?

OpenStudy (anonymous):

\[\cos (\sin^{-1} \frac{ 2 }{ 3 } + \tan^{-1} \frac{ 3 }{ 2 })\] I need to find the exact value

OpenStudy (freckles):

let a=arcsin(2/3) and let b=arctan(3/2) so sin(a)=2/3 and tan(b)=3/2 build two right triangles using that info now to expand cos(a+b) and use those triangles to evaluate cos(a+b) we will need the identity: cos(a+b)=cos(a)cos(b)-sin(a)sin(b)

OpenStudy (anonymous):

I dont understand the first part you wrote but I get the formula

OpenStudy (freckles):

hmmm... do you mean you don't know what the arc part is ? basically I said let \[a=\sin^{-1}(\frac{2}{3}) \text{ and } b=\tan^{-1} (\frac{3}{2})\]

OpenStudy (anonymous):

Oh okay I see it now

OpenStudy (anonymous):

would I leave the cos from the formula in when I input the alpha

OpenStudy (freckles):

what? not sure what you are saying have you found cos(a) given sin(a) and have you found cos(b) and sin(b) given tan(b)?

OpenStudy (freckles):

en example: